# Measurement-free code-switching for low overhead quantum computation using permutation invariant codes

### v2

@@ Appendix D Fidelity measure for Hadamard gate process
Appendix E Non-linear GPGs for switching between an even-odd code and a non even-odd code
@@ E.1 Implementing the non-linear GPG
E.2 Implementing the $\rm{C}_{\textsf{B}}{\rm X}_{\textsf{A}}$ gate
@@
E.3 Implementing the $\rm{C}_{\textsf{A}}{\rm X}_{\textsf{B}}$ gate
Second, consider the gate $\rm{C}_{\textsf{A}}{\rm X}_{\textsf{B}}$ which has the stabiliser code as the control and the PI code as target. We can use the same procedure as in Sec. E.2, with the roles of $A$ and $B$ reversed, except we choose $g\tau=\gamma=\pi$. In this case the action angle in phase space will be $\chi=4|\alpha|^{2}$ and the geometric phase gate will have the parameters $\phi=\pi$, $\theta=\pi$, and $\alpha=\beta=\sqrt{\pi}/4$. This will achieve the operation $-i\bar{Y}_{B}$ if the control is in $|1_{A}\rangle$, and acts trivially if the control is in $|0_{A}\rangle$. In summary,

### v3

@@ Appendix F Fidelity measure for Hadamard gate process
Appendix G Non-linear GPGs for switching between an even-odd code and a non even-odd code
@@ G.1 Implementing the non-linear GPG
G.2 Implementing the $\rm{C}_{\textsf{B}}{\rm X}_{\textsf{A}}$ gate
@@
G.3 Implementing the $\rm{C}_{\textsf{A}}{\rm X}_{\textsf{B}}$ gate
Second, consider the gate $\rm{C}_{\textsf{A}}{\rm X}_{\textsf{B}}$ which has the stabiliser code as the control and the PI code as target. We can use the same procedure as in Sec. G.2, with the roles of $A$ and $B$ reversed, except we choose $g\tau=\gamma=\pi$. In this case the action angle in phase space will be $\chi=4|\alpha|^{2}$ and the geometric phase gate will have the parameters $\phi=\pi$, $\theta=\pi$, and $\alpha=\beta=\sqrt{\pi}/4$. This will achieve the operation $-i\bar{Y}_{B}$ if the control is in $|1_{A}\rangle$, and acts trivially if the control is in $|0_{A}\rangle$. In summary,